# Neutrino reheating predictions with non-thermal leptogenesis

### v1

@@ 4.1 Example 1: A polynomial type potential
The results are shown in figures 1 and 2, where we show various quantities as a function of the spectral index calculated from the $N_{\mathrm{RH}}$-$N_{k}$ relation with $N_{\mathrm{RH}}$ calculated n the neutrino reheating scenario $\Gamma_{N}\ll\Gamma_{\phi}$ (18) (19). There is only one line in $N_{k}$ since it is only a function of $n_{s}$ without dependence on $M_{N}$ (28). In figure 1, a larger $M_{N}$ corresponds to a smaller $N_{\mathrm{RH}}$, $\omega_{\mathrm{RH}}$ and a higher $T_{\mathrm{RH}}$. $T_{\mathrm{RH}}$ is quite sensitive to $n_{s}$, spanning several orders of magnitude, which is also observed in ref. [11]. As the lines approach $n_{s}=0.9652$, $N_{\mathrm{RH}}$ approaches zero.
In figure 2, we show the prediction on $Y_{\mathrm{B}}$ in the $n_{s}$-$r$ plane. Since $r$ in (27) is only a function of $n_{s}$, there is only one line for $r$. $Y_{\mathrm{B}}$ (dotted red line) is degenerate in $M_{N}$. Because $M_{N}$ and $T_{\mathrm{RH}}$ show up proportionally in the $Y_{\mathrm{B}}$ expression in (12) and $N_{\mathrm{RH}}$ expression through (19). The correct amount of $Y_{\mathrm{B}}$ is marked by the red point in the figure, which corresponds to a $n_{s}$ value lying within the $1\sigma$ region of the combined Planck 2018 and BICEP/Keck results [42, 7]. As $Y_{\mathrm{B}}$ is quite sensitive to $n_{s}$, $Y_{\mathrm{B}}$ constrains the parameter space tightly around $n_{s}=0.963$ with a corresponding $T_{\mathrm{RH}}=3.25\times 10^{9}$ GeV and $r=0.146$. This tensor-to-scalar ratio is in tension with the combined Planck 2018 and BICEP/Keck results.
@@ 4.2 Example 2: Starobinsky model
$V=\Lambda^{4}\left(1-e^{-\frac{2}{3}\frac{\phi}{M_{\mathrm{P}}}}\right)^{2}\,.$ (31)
@@
$N_{k}=\frac{1}{M_{\mathrm{P}}^{2}}\int^{\phi_{k}}_{\phi_{\mathrm{end}}}\frac{M_{\mathrm{P}}}{2}\sqrt{\frac{3}{2}}\left(e^{-\frac{2}{3}\frac{\phi}{M_{\mathrm{P}}}}-1\right)\mathrm{d}\phi\simeq\frac{3}{4}e^{\frac{2}{3}\frac{\phi_{k}}{M_{\mathrm{P}}}}\,$ (32)
@@
$\phi_{k}=\sqrt{\frac{3}{2}}M_{\mathrm{P}}\mathrm{ln}\frac{4}{3}N_{k}\,,$ (33)
@@
$N_{k}=2/(1-n_{s})\,,$ (38)
substituting it back to eqs.(34) (35) (37) allow us to express all these quantities in terms of $n_{s}$. The other interested quantities are calculated in the following
@@
$V_{\mathrm{end}}$ $=\left(\frac{\frac{1}{\frac{\sqrt{3}}{2}+1}}{1-\frac{3}{8}(1-n_{s})}\right)^{2}\frac{9}{2}\pi^{2}M_{\mathrm{P}}^{4}A_{s}(1-n_{s})^{2}\,.$ (40)
@@
where $\Lambda\simeq 8\times 10^{15}$ GeV is fixed by the amplitude of the CMB anisotropies, leading to $M_{\phi}=3\times 10^{13}$ GeV.
@@
In this scenario, $N_{\mathrm{RH}}$ calculated from (18) and (19) depends on four parameters:
$M_{\phi},~{}M_{N},~{}T_{\mathrm{RH}},~{}n_{s}.$
We also choose to work with $M_{N}=10^{11},10^{9}$ GeV. Then the $N_{\mathrm{RH}}$-$N_{k}$ relation allows us to determine $T_{\mathrm{RH}}$ versus $n_{s}$. From (21) we can find an upper limit for $n_{s}$. We find a lower limit for $n_{s}$ from the requirement $N_{k}\geq 29$ following the arguments in [11].
Figure 3: Predictions in Starobinsky model with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The solid line corresponds to $M_{N}=10^{11}$ GeV and the dotted line is $M_{N}=10^{9}$ GeV. Other shadings are the same as in figure 1.
Figure 4: Predictions in Starobinsky model with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The solid blue line is the prediction on the tensor-to-scalar ratio, while the dotted red line is the prediction on the baryon asymmetry. Other shadings are the same as in figure 2.
We present the numerical results in figure 3, where various quantities are shown as functions of the spectral index. As in the polynomial potential, we find a larger $M_{N}$ corresponds to a smaller $N_{\mathrm{RH}}$, $\omega_{\mathrm{RH}}$ and a higher $T_{\mathrm{RH}}$. Planck $1\sigma$ region of $n_{s}$ corresponds to $N_{\mathrm{RH}}$ up to $28$ and $T_{\mathrm{RH}}$ in $[10^{7},10^{13}]$ GeV. In figure 4, we see that $Y_{\mathrm{B}}$ changes rapidly versus $n_{s}$ as a result of the dramatic change of $T_{\mathrm{RH}}$. The two lines in figure 4 should be read separately. The combined Planck $+$ BICEP/Keck contours constrain $r$, while the dashed light blue lines constrain $Y_{\mathrm{B}}$. We find a parameter space compatible with all observations, which is shown in table 1, where we also show $N_{k}$ and $r$ for reference although they have no dependence on $M_{N}$.
Table 1: Parameter space satisfying constraints from both the inflation and the baryon asymmetry in the Starobinsky model with $\Gamma_{N}\ll\Gamma_{\phi}$ scenario of neutrino reheating.
$M_{N}$ [GeV] $N_{\mathrm{RH}}$ $T_{\mathrm{RH}}$ [GeV] $\omega_{\mathrm{RH}}$ $N_{k}$ r
$10^{11}$ $22.63$ $3.49\times 10^{7}$ $0.074$ $51.28$ $0.0046$
$10^{9}$ $27.23$ $3.49\times 10^{5}$ $0.118$ $51.28$ $0.0046$
@@ 4.3 Example 3: Coleman-Weinberg inflation
$V=A\phi^{4}\left[\mathrm{ln}\left(\frac{\phi}{v}\right)-\frac{1}{4}\right]+\frac{Av^{4}}{4}\,$ (42)
where $v$ is the inflaton VEV at the minimum, $V(0)\equiv V_{0}=Av^{4}/4$ is the vacuum energy at origin. The slow-roll parameters are
@@
where Ei is the exponential integral function. Although the expressions may be somewhat lengthy, we can still proceed with our procedure numerically. With $\phi_{\mathrm{end}}$ calculated from eq. (43), we can determine $\phi_{k}$ as a function of $N_{k}$ with eq. (45). Hence, the functions of $\phi_{k}$ are all expressed as functions of $N_{k}$. We find $N_{k}(n_{s})$ by converting $n_{s}(N_{k})$. This facilitates us to express $H_{k}=\pi M_{\mathrm{P}}\sqrt{8A_{s}\epsilon_{k}}$ as $H_{k}(n_{s})$. For $V_{\mathrm{end}}$, we use its relation with $V_{k}$ and consequently $H_{k}$ to express it as $V_{\mathrm{end}}(n_{s})$.
The amplitude of the scalar perturbation allows us to fix $V_{0}$ for given $v$. We choose $v=30M_{\mathrm{P}}$ and $v=50M_{\mathrm{P}}$ as benchmark values, which roughly agree with inflationary observations for certain $T_{\mathrm{RH}}$ [42], and the corresponding inflaton masses are $1.78\times 10^{13}$ GeV and $1.79\times 10^{13}$ GeV [^4] In ref. [1], we use $v=22.1M_{\mathrm{P}}$ and $A=2.41\times 10^{-14}$, which agree with observation for $N_{k}=60$. Since we are interested in predictions of neutrino reheating, we fix $v$ and use scalar perturbation amplitude to fix $V_{0}$ this time. $A$ can be calculated once we know $v$ and $V_{0}$. .
@@
In this scenario $N_{\mathrm{RH}}=0$. Following the $N_{\mathrm{RH}}$-$N_{k}$ relation in (10), we find a prediction of $n_{s}=0.9652$ for $v=30M_{\mathrm{P}}$ and $n_{s}=0.9667$ for $v=50M_{\mathrm{P}}$, both are within the current $1\sigma$ region. $N_{\mathrm{RH}}=0$ also allows us to equate $\rho_{\mathrm{end}}$ with $\rho_{\mathrm{RH}}$. Hence, we find a relation of $T_{\mathrm{RH}}$ with $n_{s}$. The predicted $n_{s}$ corresponds to $T_{\mathrm{RH}}=2.92\times 10^{15}$ GeV for $v=30M_{\mathrm{P}}$ and $T_{\mathrm{RH}}=2.90\times 10^{15}$ GeV for $v=50M_{\mathrm{P}}$. The generated baryon asymmetry are $8.75\times 10^{-5}$ ($v=30M_{\mathrm{P}}$) and $8.63\times 10^{-5}$ ($v=50M_{\mathrm{P}}$), both are several orders larger than needed, leaving much space for model building.
@@
In this scenario, $N_{\mathrm{RH}}$ calculated from (18) and (19) depends on five parameters:
$M_{\phi},~{}M_{N},~{}T_{\mathrm{RH}},~{}n_{s},~{}v\,$
We also choose to work with $v=30M_{\mathrm{P}},50M_{\mathrm{P}}$ and $M_{N}=10^{11},10^{9}$ GeV. Then the $N_{\mathrm{RH}}$-$N_{k}$ relation allows us to determine $T_{\mathrm{RH}}$ versus $n_{s}$. From (21) we can find an upper limit for $n_{s}$. Following the arguments in [11], we find a lower limit for $n_{s}$ from the requirement $N_{k}\geq 29$. Note that for different $v$, the upper/lower limits for $n_{s}$ are different.
@@
Figure 5: Predictions in Coleman-Weinberg inflation with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. Other shadings are the same as in the figure 1.
Figure 6: Predictions in Coleman-Weinberg inflation with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The solid lines are predictions on the tensor-to-scalar ratio, while the dotted lines are predictions on the baryon asymmetry. The blue lines correspond to $v=30\mathrm{M}_{\mathrm{P}}$ and the red lines are $v=50\mathrm{M}_{\mathrm{P}}$. Other shadings are the same as in the figure 2.
@@
We present the numerical results in figure 5 and figure 6, where various quantities are shown as a function of the spectral index. In figure 5, we see that all lines overlap some with the observational region. Lager $v$ corresponds to a larger $N_{\mathrm{RH}}$, a lower $T_{\mathrm{RH}}$, and a smaller $N_{k}$ and $\omega_{\mathrm{RH}}$. $T_{\mathrm{RH}}$ grows with $n_{s}$ rapidly, spanning several orders of magnitudes. Given the same $v$, a larger $M_{N}$ leads to a higher $T_{\mathrm{RH}}$ and a smaller $N_{\mathrm{RH}}$ and $\omega_{\mathrm{RH}}$.
In figure 6, we show the predictions of $Y_{\mathrm{B}}$ in the $n_{s}$ $r$ plane. The predictions for $Y_{\mathrm{B}}$ and $r$ are degenerate in $M_{N}$, so there are only two lines (corresponding to $v=30M_{\mathrm{P}},50M_{\mathrm{P}}$) for each quantity. The solid lines are predictions on $r$, showing that $v=30M_{\mathrm{P}}$ is allowed while $v=50M_{\mathrm{P}}$ is not. The dotted lines are predictions for the baryon asymmetry, for which both $v$ predict a correct value within $n_{s}$ $1\sigma$ region. Combining the results, we find $v=30M_{\mathrm{P}}$ is consistent with both the $(n_{s},r)$ and $Y_{\mathrm{B}}$ constraints, while $v=50M_{\mathrm{P}}$ gives a too large $r$.

### v3

@@ 4.1 Example 1: A polynomial type potential
We choose two benchmark values of the RHN mass to show parameter correlations for interested quantities. The $N_{\mathrm{RH}}$-$N_{k}$ relation allows to determine the parameter space in terms of $T_{\mathrm{RH}}$ and $n_{s}$. Using eq. (14), we can establish a relationship between the baryon asymmetry and the spectral index. The results are shown in figure 1, where we present various quantities as functions of the spectral index calculated from the $N_{\mathrm{RH}}$-$N_{k}$ relation with $N_{\mathrm{RH}}$ calculated in the neutrino reheating scenario $\Gamma_{N}\ll\Gamma_{\phi}$. There is only one line in $N_{k}$ since it is only a function of $n_{s}$ without dependence on $M_{N}$ (28). In figure 1, a larger $M_{N}$ corresponds to a smaller $N_{\mathrm{RH}}$, $\omega_{\mathrm{RH}}$ and a higher $T_{\mathrm{RH}}$. $T_{\mathrm{RH}}$ is quite sensitive to $n_{s}$, spanning several orders of magnitude, which is also observed in ref. [11]. As the lines approach $n_{s}=0.9652$, $N_{\mathrm{RH}}$ approaches zero.
To explore the full parameter space, we notice that both $M_{N}$ and $T_{\mathrm{RH}}$ show up proportionally in the expression for $Y_{\mathrm{B}}$ in (12) and in $N_{\mathrm{RH}}$ via (19). The observed $Y_{\mathrm{B}}$ constrains only the ratio of $T_{\mathrm{RH}}/M_{N}$. In the upper-left region of the $T_{\mathrm{RH}}$-$M_{N}$ plot in the left panel of figure 2 (above the red line), the parameter space can produce a sufficiently large baryon asymmetry. Regarding the spectral index dependence, the plot in the right panel of figure 2 shows that a viable $Y_{\mathrm{B}}$ favors $n_{s}>0.9625$. Additionally, $Y_{\mathrm{B}}$ increases sharply with growing $n_{s}$, spanning several tens of orders of magnitude. The predicted tensor-to-scalar ratio is shown in figure 3, where the $n_{s}$ values preferred by successful baryogenesis are highlighted by the solid line. These results demonstrate that incorporating neutrino reheating—along with the requirement of sufficient baryon asymmetry—further restricts the parameter space.
@@ 4.2 Example 2: Starobinsky model
$V=\Lambda^{4}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{\mathrm{P}}}}\right)^{2}\,.$ (31)
@@
$N_{k}=\frac{1}{M_{\mathrm{P}}^{2}}\int^{\phi_{k}}_{\phi_{\mathrm{end}}}\frac{M_{\mathrm{P}}}{2}\sqrt{\frac{3}{2}}\left(e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{\mathrm{P}}}}-1\right)\mathrm{d}\phi\simeq\frac{3}{4}e^{\sqrt{\frac{2}{3}}\frac{\phi_{k}}{M_{\mathrm{P}}}}\,$ (32)
@@
$\phi_{k}=\sqrt{\frac{3}{2}}M_{\mathrm{P}}\mathrm{ln}\left(\frac{4}{3}N_{k}\right)\,,$ (33)
@@
$N_{k}=\frac{2}{1-n_{s}}\,,$ (38)
substituting it back to eqs.(34) (35) (37) allow us to express all these quantities in terms of $n_{s}$. With the relation in eq. (38), the other interested quantities can be expressed as a function of $n_{s}$ as
@@
$V_{\mathrm{end}}$ $=\frac{9}{2}\pi^{2}M_{\mathrm{P}}^{4}A_{s}(1-n_{s})^{2}\left(\frac{\frac{1}{\frac{\sqrt{3}}{2}+1}}{1-\frac{3}{8}(1-n_{s})}\right)^{2}\,.$ (40)
@@
where $\Lambda\simeq 8\times 10^{15}$ GeV is fixed by the amplitude of the CMB anisotropies, yielding $M_{\phi}=3\times 10^{13}$ GeV.
@@
In this scenario, $N_{\mathrm{RH}}$ calculated from (18) and (19) depends on four parameters: $M_{\phi}$, $M_{N}$, $T_{\mathrm{RH}}$, $n_{s}$. Since $M_{\phi}=3\times 10^{13}$ GeV, we are left with three variables:
$M_{N},~{}T_{\mathrm{RH}},~{}n_{s}\,.$
From (21) we can find an upper limit for $n_{s}$. We find a lower limit for $n_{s}$ from the requirement $N_{k}\geq 29$ following the arguments in [11].
Figure 4: Predictions in Starobinsky model with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The solid line corresponds to $M_{N}=10^{11}$ GeV and the dotted line is $M_{N}=10^{9}$ GeV. Other shadings are the same as in figure 1.
Figure 5: The full parameter space in Starobinsky model with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The shaded regions follow the same conventions as in figure 2.
Figure 6: Predictions for the Starobinsky model with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The red dotted line shows the predicted tensor-to-scalar ratio, while the solid portion corresponds to the range of $n_{s}$ values that yield a sufficiently large baryon asymmetry. Shadings are the same as in figure 3.
We first choose to work with $M_{N}=10^{11},10^{9}$ GeV. Then the $N_{\mathrm{RH}}$-$N_{k}$ relation allows us to determine $T_{\mathrm{RH}}$ versus $n_{s}$. The numerical results are shown in figure 4, where various quantities are displayed as functions of the spectral index. As in the polynomial potential, we find that a larger $M_{N}$ corresponds to a smaller $N_{\mathrm{RH}}$, a smaller $\omega_{\mathrm{RH}}$, and a higher $T_{\mathrm{RH}}$. Planck $1\sigma$ region of $n_{s}$ corresponds to $N_{\mathrm{RH}}$ up to $28$ and $T_{\mathrm{RH}}$ in $[10^{7},10^{13}]$ GeV.
The full parameter space is shown in figures 5 and 6. The left panel of figure 5 differs from that of figure 2 only in the upper exclusion region, due to different inflaton masses. In the right panel of figure 5, we observe that $Y_{\mathrm{B}}$ changes rapidly with $n_{s}$. A sufficiently large enough $Y_{\mathrm{B}}$ favors $n_{s}>0.961$. When projected onto the $n_{s}$-$r$ plane in figure 6, it becomes evident that including $Y_{\mathrm{B}}$ further constrains the parameter space (reducing the viable region from the entire dotted line to just the solid portion). Although $Y_{\mathrm{B}}$ prefers larger $n_{s}$ values - consistent with the latest ACT data release - achieving full agreement with observations remains challenging. The upper limit on $n_{s}$ prevents exploration of higher $n_{s}$ region.
The Starobinsky model in the Jordan frame adds the $R^{2}$ term to the Einstein-Hilbert action ($R$ is the Ricci scalar). After conformal transforming to the Einstein frame, it leads to inflaton coupling to all matter fields due to universal gravitational interactions, which is dubbed “universal reheating” [58] and is discussed in explaining BAU through leptogenesis in refs. [59, 60].
@@ 4.3 Example 3: Coleman-Weinberg inflation
$V=A\phi^{4}\left[\mathrm{ln}\left(\frac{\phi}{v}\right)-\frac{1}{4}\right]+\frac{1}{4}Av^{4}\,$ (42)
where $v$ is the inflaton vacuum expectation value at the minimum, $V(0)\equiv V_{0}=Av^{4}/4$ is the vacuum energy at origin. The amplitude of the scalar perturbation allows us to fix $V_{0}$ for given $v$. The slow-roll parameters are
@@
where Ei is the exponential integral function. Although the expressions may be somewhat lengthy, we can still proceed with our procedure numerically. With $\phi_{\mathrm{end}}$ calculated from eq. (43), we can determine $\phi_{k}$ as a function of $N_{k}$ with eq. (45). Hence, the functions of $\phi_{k}$ are all expressed as functions of $N_{k}$. We find $N_{k}(n_{s})$ by converting $n_{s}(N_{k})$. This facilitates us to express $H_{k}=\pi M_{\mathrm{P}}\sqrt{8A_{s}\epsilon_{k}}$ as $H_{k}(n_{s})$. For $V_{\mathrm{end}}$, we use its relation with $V_{k}$ and $H_{k}$ to express it as $V_{\mathrm{end}}(n_{s})$.
@@
In this scenario $N_{\mathrm{RH}}=0$. Following the $N_{\mathrm{RH}}$-$N_{k}$ relation in (10), we find a prediction of $n_{s}$ for given $v$. $N_{\mathrm{RH}}=0$ also allows us to equate $\rho_{\mathrm{end}}$ with $\rho_{\mathrm{RH}}$. Hence, we find a relation of $T_{\mathrm{RH}}$ with $n_{s}$. The predictions correspond to the ending points in figure 9. The intersection with the Planck-BK18 $2\sigma$ region leads to the following estimation of the parameter space:
$v\in[13,47]M_{\mathrm{P}}$
$T_{\mathrm{RH}}\in[2.47,2.91]\times 10^{15}\mathrm{GeV}$
$M_{\phi}\in[1.40,1.71]\times 10^{13}\mathrm{GeV}\,.$
The predicted baryon asymmetry are from $9.05\times 10^{-5}$ to $9.39\times 10^{-5}$, several orders larger than needed, leaving much space for model building.
@@
[…]
$M_{N},~{}T_{\mathrm{RH}},~{}v,~{}n_{s}\,.$
@@
Figure 7: Predictions in Coleman-Weinberg inflation with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. Other shadings are the same as in the figure 1.
@@
To illustrate the correlations among the relevant quantities, we first consider benchmark values of $v=30M_{\mathrm{P}},50M_{\mathrm{P}}$ and $M_{N}=10^{11},10^{9}$ GeV. Using the $N_{\mathrm{RH}}$-$N_{k}$ relation, we then determine $T_{\mathrm{RH}}$ as a function of $n_{s}$. The results are shown in figure 7, where various quantities are shown as functions of the spectral index. In figure 7, we see that all lines overlap some with the observational region. Increasing the parameter $v$ leads to an extended reheating duration $N_{\mathrm{RH}}$, a reduced reheating temperature $T_{\mathrm{RH}}$, and smaller values of both $N_{k}$ and $\omega_{\mathrm{RH}}$. $T_{\mathrm{RH}}$ grows with $n_{s}$ rapidly, spanning several orders of magnitude. For a fixed $v$, a larger $M_{N}$ corresponds to a higher $T_{\mathrm{RH}}$, as well as smaller values of $N_{\mathrm{RH}}$ and $\omega_{\mathrm{RH}}$.
Figure 8: The full parameter space in Coleman-Weinberg inflation with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. In the left panel, horizontal and vertical dashed lines indicate the upper limits of $T_{\mathrm{RH}}$ and $M_{N}$ respectively, determined by different values of $v$. Other shadings are the same as in the figure 2.
Figure 9: Predictions in Coleman-Weinberg inflation with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The red dotted lines are the predicted tensor-to-scalar ratio. The continuous parts of the red lines indicate $n_{s}$ values generating sufficiently large $Y_{\mathrm{B}}$. Shadings are the same as in figure 3.
The full parameter space is shown in figures 8 and 9. In the left panel of figure 8, the gray shaded region at the top is excluded by the requirement of $N_{\mathrm{I}}>0$, which imposes an upper bound on $T_{\mathrm{RH}}$ that varies with $v$. The dotted lines mark five $v$ values from $10M_{\mathrm{P}}$ to $50M_{\mathrm{P}}$, from bottom to top is the increasing order $v$. The upper three lines are nearly indistinguishable due to their close spacing. The same happens for the five vertical dotted lines, which show the kinetic upper limit of $M_{N}$. The viable parameter space that produces sufficient baryon asymmetry $Y_{\mathrm{B}}$ lies above the red line and below the dotted $T_{\mathrm{RH}}$ limit. The $M_{N}$ upper bounds are less restrictive than the $Y_{\mathrm{B}}$ constraint.
The right panel of figure 8 displays the dependence of the baryon asymmetry $Y_{\mathrm{B}}$ on the spectral index $n_{s}$ with different values of $v$. We observe that $Y_{\mathrm{B}}$ increases sharply with $n_{s}$. The lower bounds of $n_{s}$, determined by the requirement $N_{k}\geq 29$, lie outside the plotted range. Using (21), we derive and plot upper limits for $n_{s}$. $v<20M_{\mathrm{P}}$ can hardly overlap with the $n_{s}$ $1\sigma$ region. A larger $v$ corresponds to a larger $n_{s}$.
Figure 9 presents our predictions for the tensor-to-scalar ratio. The dotted curves show the predicted $r$, with their solid segments indicating the $n_{s}$ range capable of producing sufficient baryon asymmetry. The plot demonstrates that including the baryon asymmetry constraint $Y_{\mathrm{B}}$ substantially reduces the allowed parameter space. For a given spectral index $n_{s}$, a larger $v$ corresponds to a larger $r$. Consequently, the allowed range of $v$ is constrained by $r$ from the top and $n_{s}$ from the bottom. While showing no overlap with ACT-inclusive datasets, our results exhibit partial agreement with Planck-BK18 data for the parameter range $v\sim[13,47]M_{\mathrm{P}}$.